# MERT: Acoustic Music Understanding Model with Large-Scale Self-supervised Training

### v3

@@ Abstract
Self-supervised learning (SSL) has recently emerged as a promising paradigm for training generalisable models on large-scale data in the fields of vision, text, and speech. Although SSL has been proven effective in speech and audio, its application to music audio has yet to be thoroughly explored. This is partially due to the distinctive challenges associated with modelling musical knowledge, particularly tonal and pitched characteristics of music. To address this research gap, we propose an acoustic Music undERstanding model with large-scale self-supervised Training (MERT), which incorporates teacher models to provide pseudo labels in the masked language modelling (MLM) style acoustic pre-training. In our exploration, we identified an effective combination of teacher models, which outperforms conventional speech and audio approaches in terms of performance. This combination includes an acoustic teacher based on Residual Vector Quantisation - Variational AutoEncoder (RVQ-VAE) and a musical teacher based on the Constant-Q Transform (CQT). Furthermore, we explore a wide range of settings to overcome the instability in acoustic language model pre-training, which allows our designed paradigm to scale from 95M to 330M parameters. Experimental results indicate that our model can generalise and perform well on 14 music understanding tasks and attain state-of-the-art (SOTA) overall scores.
@@ 1 Introduction
Pre-trained language models (PLMs) can learn generalisable representations of data without human annotated labels in a self-supervised learning (SSL) style, leading to remarkable performance improvement in natural language processing and related fields (Brown et al.,, 2020; Fang et al.,, 2022; Chen et al., 2021a, ). Music is widely recognised as a special language that can be used to communicate across different cultures  (Mehr et al.,, 2019). The internal similarity between music and language as a communication interface lays a promising foundation for adapting PLM-based methods to model music sequences. We argue that the benefit is twofold. First, PLMs can potentially pave the way to unify the modelling of a wide range of music understanding, or the so-called Music Information Retrieval (MIR) tasks, including but not limited to music tagging, beat tracking, music transcription, source separation, etc., so that different tasks no longer need task-specific models or features. Second, releasing a PLM for acoustic music understanding allows the redistribution of the musical knowledge rather than the data itself, which saves the costs of manual annotation and copyright law restrictions.
Unfortunately, we are yet to see a general-purpose and cost-effective open-source PLM on acoustic music understanding. Most existing studies are designed to solely address music tagging problems (Pons and Serra,, 2019; Spijkervet and Burgoyne,, 2021; McCallum et al.,, 2022; Huang et al.,, 2022; Zhu et al.,, 2021; Zhao and Guo,, 2021), and many of them do not provide open-source code bases or checkpoints for further evaluation. A promising model is JukeMIR (Castellon et al.,, 2021), which is based on Jukebox (Dhariwal et al.,, 2020) and provides a comprehensive evaluation on MIR downstream tasks. However, this foundation model uses cumbersome hierarchical auto-regressive transformer decoders containing billions of parameters to model music audio, leading to significant inefficiency for conducting general music understanding tasks (e.g., it takes weeks to inference on datasets like MTG (Bogdanov et al.,, 2019) with a consumer-grade 3090 GPU).
@@
The aforementioned research gap has urged us to design and open-source a generalisable and computationally affordable pre-trained acoustic music model. In this paper, we propose an acoustic Music undERstanding model with large-scale self-supervised Training (MERT). MERT inherits a speech SSL paradigm, employing teacher models to generate pseudo targets for sequential audio clips. Specifically, to capture the distinctive pitched and tonal characteristics in music, MERT incorporates a multi-task paradigm to balance the acoustic and musical representation learning as demonstrated in Fig. 1. In the proposed design, an Residual Vector Quantisation - Variational Autoencoder (RVQ-VAE) (Défossez et al.,, 2022) is used as the acoustic teacher to provide discretised acoustic-level summarisation of the music signal. The Constant-Q Transformation (CQT) (Brown,, 1991) model is further introduced as the music teacher for capturing pitch and harmonic inductive bias. Regarding the context dependencies and music hierarchies, as indicated in Borsos et al., (2022), we leave the task of modelling high-level and abstract patterns to the stacked layers of self-attentions in the transformer.
@@
  • conducting a broad range of analysis based on ablation study of the proposed MERT pre-training paradigm;
@@ 3.1 Pre-training with MLM
where $\mathbf{1}_{[L]\setminus M}(x)$ denotes the indicator function, that is, $\mathbf{1}_{[L]\setminus M}(x)=1$ if and only if $x$ is outside the masked indices set $M$. The pseudo-label that needs to be learned is typically $y_{i}^{\prime}=x_{i}-x_{i}^{\prime}$, i.e., the masked data. However, reconstructing masked data $y^{\prime}$ for raw audio tasks as pseudo-label is hard to train. HuBERT (Vaswani et al.,, 2017; Hsu et al.,, 2021) uses a dimension-reduced feature $z^{\prime}$ derived from $y^{\prime}$ with phonetic acoustic information, which forms the design basis of our pre-training strategy.
@@ 3.2 Modelling Acoustic Information
The MFCC features are only good at modelling acoustic and timbre information for single-pitch signals, and therefore, the clustering results do not provide much timbre information in music recording. We proposed two potential approaches as the teacher on acoustic information: one based on traditional features, and the other based on deep learning.
The first method uses k-means on the log-Mel spectrum and Chroma features for timbre and harmonic acoustic information, respectively. In the case of music representation, each frame contains more information compared to speech, necessitating a larger number of classes for k-means clustering. The complexity of the k-means algorithm is linear with the number of centroids (clustering centres), leading to a time-consuming k-means for the music feature. To tackle this problem, we employ 300-means for the log-Mel spectrum with dimension 229, and 200-means for Chroma features with dimension 264, resulting in a total of 60,000 classes (200 centroids for Chroma features multiplied by 300 centroids for the log-Mel spectrum). Despite the increased number of classes, the computational complexity remains comparable to that of HuBERT. The disadvantage of k-means is that it is difficult to scale up to a larger number of classes and larger datasets, and the results are sensitive to initialisation.
@@ 4.1 Evaluation Protocol
Following Castellon et al., (2021); Yang et al., (2021), we restrict the testing protocol with probing rather than fine-tuning, i.e. freezing the backbone pre-trained models as deep feature extractor and only train a simple downstream structure, typically a multilayer perceptron (MLP) for frame-level tasks. For a fair comparison, we also limit the space for hyper-parameters searching. For full details please refer to Appendix B.2.
@@ 4.3 Implementation Details
We deploy the proposed SSL architecture in the training of various model sizes with matched scales of data. We mined 160K hours of music recordings from the Internet to build a large-scale music dataset. Accordingly, the base size models (95M) are trained with a 1K hours subset whereas the whole dataset is used for the large model (330M). Specifically, we provide a special edition of the base model, MERT-95M-public, that is trained on a totally publicly available music dataset, music4all (Santana et al.,, 2020), with a data size of 910 hours. In the context of self-attention, the computational complexity scales quadratically with the sequence length. Therefore, when dealing with limited computational resources, there exists a trade-off between the batch size and the sequence length. In our preliminary experiments, we have observed that increasing the batch size provides greater performance improvements compared to extending the context length. To allow a larger batch size under the computational limitation, we adopt a strategy of randomly truncating audio clips into 5-second segments following Ma et al., (2023). This duration roughly corresponds to a 2-bar context in music. It is worth noting that our model utilises a convolutional relative positional embedding, similar to the approach introduced by Baevski et al., (2020) in Wav2Vec, enabling it to operate effectively in longer contexts, if required. The effective batch sizes and learning rates for the base model and large model are set to $1.5$ and $5.5$ hours, and their learning rates are set to $5\mathrm{e}{-4}$, $1.5\mathrm{e}{-3}$, respectively. Pre-training is carried out with the fairseq [^1] https://github.com/facebookresearch/fairseq framework. Models are trained with 64 A100-40GB GPUs with fp16. We also implement a data augmentation of randomly adding short segments to improve the representation robustness, and describe the details in Appendix A.1
@@
Table 1: Experimental Performances of MERT and Baselines on Downstream Tasks (1/2). The baselines are grouped by supervised and unsupervised pre-training paradigms. The superscripts denote the category of the acoustic teacher used by MERT models. “public” refers to the MERT model trained with only open-source dataset. Results with star* are claimed in the references.
@@
Table 2: Experimental Performances of MERT and Baselines on Downstream Tasks (2/2). Average scores across task are calculated on the SOTA results and models applicable to all the tasks.
@@ 5.1 Performance & Efficiency of MERT Models
Table 3: Evaluation Results from Models Trained with Different Teacher Settings. Models labeled with ${}^{\vartriangle 2}$ and ${}^{\blacktriangle 2}$ suggest that the K-means teachers are trained with the features from ${}^{\vartriangle 1}$ and ${}^{\blacktriangle 1}$ models. All the listed models are sized in (95M) and not augmented with the in-batch noise mixture.
@@
[…]
K-meansLogmel+Chroma 300 + 200 ${}^{\blacktriangle 1}$ 90.5 37.6 55.1 75.2 40.1 68.2 62.1
K-meansMFCC ${}^{\vartriangle 2}$ 90.4 37.5 16.1 68.3 43.9 67.7 51.0
[…]
@@ 6 Conclusion
Limitation and Future Work
@@ Appendix D Ethics
We have taken great care to ensure that our research adheres to ethical principles and guidelines in the codes of conduct. Specifically, we have not used inappropriate user information in the experiments. The audios are collected from open-access streaming services and open-source datasets, where the quality of the audios vary among 16KHz, 24KHz, and 48KHz. The audios would not be hosted and distributed. We believe that our work has the potential to contribute to positive social and scientific outcomes regarding the research of automatic music understanding.

### v5

@@ Abstract
Self-supervised learning (SSL) has recently emerged as a promising paradigm for training generalisable models on large-scale data in the fields of vision, text, and speech. Although SSL has been proven effective in speech and audio, its application to music audio has yet to be thoroughly explored. This is partially due to the distinctive challenges associated with modelling musical knowledge, particularly tonal and pitched characteristics of music. To address this research gap, we propose an acoustic Music undERstanding model with large-scale self-supervised Training (MERT), which incorporates teacher models to provide pseudo labels in the masked language modelling (MLM) style acoustic pre-training. In our exploration, we identified an effective combination of teacher models, which outperforms conventional speech and audio approaches in terms of performance. This combination includes an acoustic teacher based on Residual Vector Quantisation - Variational AutoEncoder (RVQ-VAE) and a musical teacher based on the Constant-Q Transform (CQT). Furthermore, we explore a wide range of settings to overcome the instability in acoustic language model pre-training, which allows our designed paradigm to scale from 95M to 330M parameters. Experimental results indicate that our model can generalise and perform well on 14 music understanding tasks and attain state-of-the-art (SOTA) overall scores. The code and models are online: https://github.com/yizhilll/MERT.
@@ 1 Introduction
Pre-trained language models (PLMs) can learn generalisable representations of data without human annotated labels in a self-supervised learning (SSL) style, leading to remarkable performance improvement in natural language processing and related fields (Brown et al.,, 2020; Fang et al.,, 2022; Chen et al., 2021a, ). Music is widely recognised as a special language that can be used to communicate across different cultures  (Mehr et al.,, 2019). The internal similarity between music and language as a communication interface lays a promising foundation for adapting PLM-based methods to model music sequences. We argue that the benefit is twofold. First, PLMs can potentially pave the way to unify the modelling of a wide range of music understanding, or the so-called Music Information Retrieval (MIR) tasks, including but not limited to music tagging, beat tracking, music transcription, and source separation, so that different tasks no longer need task-specific models or features. Second, releasing a PLM for acoustic music understanding allows the redistribution of the musical knowledge rather than the data itself, which saves the costs of manual annotation and copyright law restrictions.
Unfortunately, we are yet to see a general-purpose and cost-effective open-source PLM on acoustic music understanding. Most existing studies are designed to solely address music tagging problems (Pons and Serra,, 2019; Spijkervet and Burgoyne,, 2021; McCallum et al.,, 2022; Huang et al.,, 2022; Zhu et al.,, 2021; Zhao and Guo,, 2021), and many of them do not provide open-source code bases or checkpoints for further evaluation. A promising model is JukeMIR (Castellon et al.,, 2021), which is based on Jukebox (Dhariwal et al.,, 2020) and provides a comprehensive evaluation on MIR downstream tasks. However, this foundation model uses cumbersome hierarchical auto-regressive transformer decoders containing billions of parameters to model music audio, leading to significant inefficiency for conducting general music understanding tasks (e.g., it takes weeks to carry out inference on datasets like MTG (Bogdanov et al.,, 2019) with a consumer-grade 3090 GPU).
@@
The aforementioned research gap has urged us to design and open-source a generalisable and computationally affordable pre-trained acoustic music model. In this paper, we propose an acoustic Music undERstanding model with large-scale self-supervised Training (MERT). MERT inherits a speech SSL paradigm, employing teacher models to generate pseudo targets for sequential audio clips. Specifically, to capture the distinctive pitched and tonal characteristics in music, MERT incorporates a multi-task paradigm to balance the acoustic and musical representation learning as demonstrated in Fig. 1. In the proposed design, a Residual Vector Quantisation - Variational Autoencoder (RVQ-VAE) (Défossez et al.,, 2022) is used as the acoustic teacher to provide discretised acoustic-level summarisation of the music signal. The Constant-Q Transformation (CQT) (Brown,, 1991) model is further introduced as the music teacher for capturing pitch and harmonic inductive bias. Regarding the context dependencies and music hierarchies, as indicated in Borsos et al., (2022), we leave the task of modelling high-level and abstract patterns to the stacked layers of self-attentions in the transformer.
@@
  • conducting a broad range of analysis based on ablation studies of the proposed MERT pre-training paradigm;
@@ 3.1 Pre-training with MLM
where $\mathbf{1}_{[L]\setminus M}(x)$ denotes the indicator function, that is, $\mathbf{1}_{[L]\setminus M}(x)=1$ if and only if $x$ is outside the masked indices set $M$. The pseudo-label that needs to be learned is typically $y_{i}^{\prime}=x_{i}-x_{i}^{\prime}$, i.e., the masked data. However, reconstructing masked data $y^{\prime}$ for raw audio tasks as pseudo-labels is hard to train. HuBERT (Vaswani et al.,, 2017; Hsu et al.,, 2021) uses a dimension-reduced feature $z^{\prime}$ derived from $y^{\prime}$ with phonetic acoustic information, which forms the design basis of our pre-training strategy.
@@ 3.2 Modelling Acoustic Information
The MFCC features are only good at modelling acoustic and timbre information for single-pitch signals, and therefore, the clustering results do not provide much timbre information in music recordings. We proposed two potential approaches as the teacher on acoustic information: one based on traditional features, and the other based on deep learning.
The first method uses k-means on the log-Mel spectrum and Chroma features for timbre and harmonic acoustic information, respectively. In the case of music representations, each frame contains more information compared to speech, necessitating a larger number of classes for k-means clustering. The complexity of the k-means algorithm is linear with the number of centroids (clustering centres), leading to a time-consuming k-means for the music feature. To tackle this problem, we employ 300-means for the log-Mel spectrum with dimension 229, and 200-means for Chroma features with dimension 264, resulting in a total of 60,000 classes (200 centroids for Chroma features multiplied by 300 centroids for the log-Mel spectrum). Despite the increased number of classes, the computational complexity remains comparable to that of HuBERT. The disadvantage of k-means is that it is difficult to scale up to a larger number of classes and larger datasets, and the results are sensitive to initialisation.
@@ 4.1 Evaluation Protocol
Following Castellon et al., (2021); Yang et al., (2021), we restrict the testing protocol with probing rather than fine-tuning, i.e. freezing the backbone pre-trained models as deep feature extractor and only train a simple downstream structure, typically a multilayer perceptron (MLP) for frame-level tasks. For a fair comparison, we also limit the space for hyper-parameter search. For full details please refer to Appendix B.2.
@@ 4.3 Implementation Details
We deploy the proposed SSL architecture in the training of various model sizes with matched scales of data. We mined 160K hours of music recordings from the Internet to build a large-scale music dataset. Accordingly, the base size models (95M) are trained with a 1K hours subset whereas the whole dataset is used for the large model (330M). Specifically, we provide a special edition of the base model, MERT-95M-public, that is trained on a totally publicly available music dataset, music4all (Santana et al.,, 2020), with a data size of 910 hours. In the context of self-attention, the computational complexity scales quadratically with the sequence length. Therefore, when dealing with limited computational resources, there exists a trade-off between the batch size and the sequence length. In our preliminary experiments, we have observed that increasing the batch size provides greater performance improvements compared to extending the context length. To allow a larger batch size under the computational limitation, we adopt a strategy of randomly truncating audio clips into 5-second segments following Ma et al., (2023). This duration roughly corresponds to a 2-bar context in music. It is worth noting that our model utilises a convolutional relative positional embedding, similar to the approach introduced by Baevski et al., (2020) in Wav2Vec, enabling it to operate effectively in longer contexts, if required. The effective batch sizes and learning rates for the base model and large model are set to $1.5$ and $5.5$ hours, and their learning rates are set to $5\mathrm{e}{-4}$, $1.5\mathrm{e}{-3}$, respectively. Pre-training is carried out with the fairseq [^1] https://github.com/facebookresearch/fairseq framework. Models are trained with 64 A100-40GB GPUs with fp16. We also implement a data augmentation of randomly adding short segments to improve the representation robustness, and describe the details in Appendix A.1.
@@
Table 1: Experimental Performance of MERT and Baselines on Downstream Tasks (1/2). The baselines are grouped by supervised and unsupervised pre-training paradigms. The superscripts denote the category of the acoustic teacher used by MERT models. “public” refers to the MERT model trained with only open-source dataset. Results with star* are claimed in the references.
@@
Table 2: Experimental Performance of MERT and Baselines on Downstream Tasks (2/2). Average scores across task are calculated on the SOTA results and models applicable to all the tasks.
@@ 5.1 Performance & Efficiency of MERT Models
Table 3: Evaluation Results from Models Trained with Different Teacher Settings. Models labeled with △2 and ▲2 suggest that the K-means teachers are trained with the features from △1 and ▲1 models. All the listed models are sized in (95M) and not augmented with the in-batch noise mixture.
@@
K-meansMFCC 2000△1 90.2 37.6 15.6 70 44.3 67.6 51.4
K-meansLogmel+Chroma 300 + 200 ▲1 90.5 37.6 55.1 75.2 40.1 68.2 62.1
K-meansMFCC 2000△2 90.4 37.5 16.1 68.3 43.9 67.7 51.0
K-meansLogmel+Chroma 500▲2 90.4 37.7 49.2 72.8 46.5 66.9 60.7
@@ 6 Conclusion
Limitations and Future Work
@@ Appendix D Ethics
We have taken great care to ensure that our research adheres to ethical principles and guidelines in the codes of conduct. Specifically, we have not used personal user information in the experiments. Audio recordings are collected from open-access streaming services and open-source datasets, where the quality of the recordings varies among 16KHz, 24KHz, and 48KHz. The audio data would not be hosted and distributed. We believe that our work has the potential to contribute to positive social and scientific outcomes regarding the research of automatic music understanding.